# Revisiting PSF models: unifying framework and high-performance implementation

### v1

@@ Abstract
Localization microscopy often relies on detailed models of point spread functions. For applications such as deconvolution or PSF engineering, accurate models for light propagation in imaging systems with high numerical aperture are required. Different models have been proposed based on 2D Fourier transforms or 1D Bessel integrals. The most precise ones combine a vectorial description of the electric field and precise aberration models. However, it may be unclear which model to choose, as there is no comprehensive comparison between the Fourier and Bessel approaches yet. Moreover, many existing libraries are written in Java (e.g. our previous PSF generator software) or MATLAB, which hinders the integration into deep learning algorithms. In this work, we start from the original Richards-Wolf integral and revisit both approaches in a systematic way. We present a unifying framework in which we prove the equivalence between the Fourier and Bessel strategies and detail a variety of correction factors applicable to both of them. Then, we provide a high-performance implementation of our theoretical framework in the form of an open-source library that is built on top of PyTorch, a popular library for deep learning. It enables us to benchmark the accuracy and computational speed of different models, thus allowing for an in-depth comparison of the existing models for the first time. We show that the Bessel strategy is optimal for axisymmetric beams while the Fourier approach can be applied to more general scenarios. Our work enables efficient PSF computation on CPU or GPU, which can then be included in simulation and optimization pipelines.
Keywords point spread function  $\cdot$ localization microscopy  $\cdot$ vectorial field propagation  $\cdot$ open-source library
@@ 1 Introduction
The point spread function (PSF), also referred to as the impulse response, is a fundamental concept that encapsulates the key features of an optical microscope. Its monitoring and assessment have been long-standing routines in microscopy, supported by the development of specific tools [1, 2, 3] and the creation of a consortium to standardize best practices [4]. Detailed characterization of the PSF is essential to design computational imaging processes, such as single-molecule localization microscopy [5] and super-resolution microscopy, including 3D deconvolution microscopy [6, 7], structured illumination microscopy (SR-SIM) [8, 9], fluctuation microscopy [10, 11], stimulated emission-depletion microscopy (STED) [12], and MINFLUX [13]. The PSF is at the heart of these localization techniques to achieve a spatial resolution beyond the optical diffraction limit.
@@
The PSF can be measured experimentally from a z-stack of tiny fluorescent beads [6, 14]. However, a theoretical PSF is often required. Here, the challenge lies in the accurate model of light propagation for imaging systems with high numerical aperture (NA). Detailed models have been proposed [15, 16, 17, 18] to (a) go beyond simple paraxial approximations, (b) take into account the vectorial nature of the electric field, and (c) include various aberration factors, such as Gibson-Lanni aberrations due to refraction at different planar interfaces [19]. Several software packages have been developed to generate precise theoretical 3D PSFs that include various features and different aberrations. Among these, the Huygens PSF software [^1] https://svi.nl/ is well-known in the microscopy community, and our Java-based ImageJ/FIJI plugin, PSF Generator [20], has also been widely used, though it is now outdated in terms of implementation and software design. Other open-source alternatives include MATLAB-based solutions [21, 3, 22] as well as Python-based tools [23, 24].
@@
These computational PSF models have been applied in various applications in fluorescence microscopy. For example, fitting a theoretical PSF to a measured one can improve accuracy [20, 25]. Detailed models enable quantification of 3D uncertainty [26] and PSF engineering takes full advantage of parametric PSF models [27, 28, 29]. They are a key component to create for example high-quality SMLM images [30, 31], including single-particle tracking [32, 33], and virtual SMLM microscopes [34, 35].
@@
Existing approaches to compute the PSF can be categorized into two classes: those based on 2D Fourier transforms [16, 3] or 1D integrals of Bessel functions [15, 36, 17, 18]. These two classes of models have different assumptions (Bessel models being more restrictive) and bring different computational trade-offs. Most of these works focus on one of the two approaches, and the relationship between them remains unclear. For instance, the Gibson-Lanni aberrations [19] or the apodization factor [15] are only applied to the Bessel case, even though they could be generalized to both. To the best of our knowledge, there has not been a systematic benchmark of the two strategies yet in terms of accuracy and computational speed.
@@
In this work, we propose a unifying framework for PSF models, in which we show the equivalence between the Fourier and Bessel approaches as different parametrizations of the same propagation integral. This enables us to generalize diverse correction factors and apply them to both models. Eventually, the choice is simplified to scalar versus vectorial models, optionally with additional correction factors. We then distribute an open-source PyTorch-based library, called psf-generator, which inherits all the functionalities of PyTorch and allows for a seamless integration with modern learning-based algorithms. We benchmark the two classes of models on CPU and GPU and show that the Bessel model is advantageous for axisymmetric PSFs while the Fourier one is more general, and thus more suitable for applications such as PSF engineering. Our benchmark can be updated regularly to follow the latest hardware and software developments.
@@ 2 Background
Electromagnetic waves in optical systems are fundamentally described by the Helmholtz wave equations yet solving these equations in full generality is computationally intractable. Various approximations and integral formulations have thus been developed. The cornerstone for precise PSF models in high-NA systems is the Richards-Wolf integral [15], which can be viewed as a vectorial extension of the Debye integral [37]. The conditions of validity for the Richards-Wolf integral have been thoroughly discussed by [38]. While alternative light propagation models exist, the Huygens-Fresnel approach has been shown to be equivalent, to some extent, for PSF calculations [39].
@@ 3 Theory
Figure 1: Geometry of the focusing of optical fields. An incident field $\mathbf{e}_{\mathrm{inc}}$ is transformed by a focusing element into a converging spherical wave $\mathbf{e}_{\infty}$. These fields are parametrized by a unit vector $\mathbf{s}$, either in Cartesian coordinates $(s_{x},s_{y},s_{z})$ or spherical coordinates $(\theta,\phi)$. The focus field $\mathbf{E}$ is parametrized by $\boldsymbol{\rho}=(x,y,z)$, rewritten here to introduce cylindrical coordinates $\rho$ and $\phi$.
@@ 3.1 The Richards-Wolf Model
As depicted in Fig. 1, the PSF is obtained by computing the propagation of light after going through a focusing element, typically a microscope objective or lens. The incident field $\mathbf{e}_{\mathrm{inc}}(\mathbf{s})$, also called pupil function, is represented by a disk with a maximal cut-off angle defined by the NA of the imaging system. The focusing element transforms the incident field into a spherical wave, $\mathbf{e}_{\infty}(\mathbf{s})$, evaluated on the Gaussian reference sphere. This corresponds to an ensemble of far fields propagating with direction $\mathbf{s}$ all converging to the focal point $O$. Our goal is to compute the focused electric field $\mathbf{E}(\boldsymbol{\rho})$ around the point $O$. Thanks to the reciprocity of light propagation, this model can also be extended to the emission of a point source to the back focal plane of a microscope objective but this is not the focus of the current study.
The model introduced by [15] is the starting point that allows us to derive all the precise PSF models described in previous works. The focal field is given by a sum of plane waves with direction $\mathbf{s}=(s_{x},s_{y},s_{z})$:
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{\Omega}\mathbf{e}_{\infty}(\mathbf{s})\exp{\left\{\mathrm{i}k\mathbf{s}\cdot\boldsymbol{\rho}\right\}}\mathrm{d}\Omega$ (1)
where $\boldsymbol{\rho}=(x,y,z)=(\rho\cos\varphi,\rho\sin\varphi,z)$ is the position vector in the focal region of the lens, $\mathbf{s}=(s_{x},s_{y},s_{z})=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$ is a unit vector describing the direction of an incoming ray, $f$ is the focal length of the lens, $k=\frac{2\pi n}{\lambda}$ is the wavenumber, $\lambda$ is the wavelength, $n$ is the refractive index of the propagation medium, and $\mathbf{e}_{\infty}(\mathbf{s})$ describes the field distribution on the Gaussian reference sphere. We integrate over set $\Omega$ of solid angles defined on a region $s_{x}^{2}+s_{y}^{2}\leq s_{\mathrm{max}}^{2}$, where $s_{\mathrm{max}}=\frac{\mathrm{NA}}{n_{i}}$ is the cut-off determined by the NA. The angle $\theta$ is defined in the immersion medium. Correction factors are typically introduced in this expression but we will first describe the different classes of models based on this simplified concise equation.
@@ 3.2 Scalar Models
As a first step, it is common to employ a scalar approximation to simplify calculations, especially in low-NA scenarios. In this case, the far field is equal to the incident field, i.e. $e_{\infty}(\mathbf{s})=e_{\mathrm{inc}}(\mathbf{s})$ and the focal field is given by:
$E(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{\Omega}e_{\infty}(s_{x},s_{y})\exp{\left\{\mathrm{i}k\mathbf{s}\cdot\boldsymbol{\rho}\right\}}\mathrm{d}\Omega.$ (2)
This expression involves a two-dimensional integral over the pupil disk. Two possible parametrizations that yield the two classes of models described previously can be employed.
The Cartesian parametrization utilizes both $s_{x}$ and $s_{y}$ coordinates with $\mathrm{d}\Omega=\mathrm{d}s_{x}\mathrm{d}s_{y}/s_{z}$, resulting in:
$E(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{s_{x}^{2}+s_{y}^{2}\leq s_{\max}^{2}}\frac{e_{\infty}(s_{x},s_{y})}{s_{z}}\exp{\left\{\mathrm{i}ks_{z}z\right\}}\exp{\left\{\mathrm{i}k(s_{x}x+s_{y}y)\right\}}\mathrm{d}s_{x}\mathrm{d}s_{y}.$ (3)
In this form, the focused field at a given transverse plane is given by the 2D inverse Fourier transform of $(e_{\infty}(s_{x},s_{y})\exp{\left\{\mathrm{i}ks_{z}z\right\}}/s_{z})$, where $s_{z}=\sqrt{1-s_{x}^{2}-s_{y}^{2}}$. Thus, the Cartesian parametrization of the Richards-Wolf integral leverages the speed and efficiency of the Fast Fourier Transform (FFT) algorithm.
Alternatively, the spherical approach parametrizes the problem with two angles $\theta\in[0,\theta_{\max}]$ (the maximum angle $\theta_{\max}$ is determined by the NA) and $\phi\in[0,2\pi]$, as depicted in Figure 1. With $\boldsymbol{\rho}=(x,y,z)=(\rho\cos\varphi,\rho\sin\varphi,z)$ and $\mathrm{d}\Omega=\sin\theta\mathrm{d}\theta\mathrm{d}\phi$, the field in the focal region can be rewritten as:
@@
$\exp{\left\{\mathrm{i}k\rho\sin\theta\cos(\phi-\varphi)\right\}}\exp{\left\{\mathrm{i}kz\cos\theta\right\}}\sin\theta.$ (4)
Eq. (3.2) can be further simplified if one assumes that the pupil function is axisymmetric (rotational invariant), i.e. $e_{\infty}(\theta,\phi)=e_{\infty}(\theta)$. In this case the integral over $\phi$ in Eq. (3.2) can be computed explicitly using the Bessel function $J_{0}$ [^2] The spherical parametrization often uses the following identities where $J_{n}$ is the Bessel function of $n$th-order of the first kind: $\int_{0}^{2\pi}\cos(n\phi)\exp{\left\{\mathrm{i}x\cos(\phi-\varphi)\right\}}\mathrm{d}\phi=2\pi(\mathrm{i})^{n}J_{n}(x)\cos(n\varphi)$ $\int_{0}^{2\pi}\sin(n\phi)\exp{\left\{\mathrm{i}x\cos(\phi-\varphi)\right\}}\mathrm{d}\phi=2\pi(\mathrm{i})^{n}J_{n}(x)\sin(n\varphi)$ :
$E(\boldsymbol{\rho})=-\mathrm{i}fk\int_{0}^{\theta_{\max}}\mathrm{d}\theta e_{\infty}(\theta)J_{0}(k\rho\sin\theta)\exp{\left\{\mathrm{i}kz\cos\theta\right\}}\sin\theta.$ (5)
Defocus is included in these models using the defocus phase factor $\exp{\left\{\mathrm{i}ks_{z}z\right\}}=\exp{\left\{\mathrm{i}kz\cos\theta\right\}}$ where $z$ is the defocus distance. This expression, also known as angular spectrum propagation [41], accurately models the propagation of electric field in a homogeneous medium.

### v2

@@ Abstract
Localization microscopy often relies on detailed models of point-spread functions. For applications such as deconvolution or PSF engineering, accurate models for light propagation in imaging systems with a high numerical aperture are required. Different models have been proposed based on 2D Fourier transforms or 1D Bessel integrals. The most precise ones combine a vectorial description of the electric field and accurate aberration models. However, it may be unclear which model to choose as there is no comprehensive comparison between the Fourier and Bessel approaches yet. Moreover, many existing libraries are written in Java (e.g., our previous PSF generator software) or MATLAB, which hinders their integration into deep learning algorithms. In this work, we start from the original Richards-Wolf integral and revisit both approaches in a systematic way. We present a unifying framework in which we prove the equivalence between the Fourier and Bessel strategies and detail a variety of correction factors applicable to both of them. Then, we provide a high-performance implementation of our theoretical framework in the form of an open-source library that is built on top of PyTorch, a popular library for deep learning. It enables us to benchmark the accuracy and computational speed of different models and allows for an in-depth comparison of the existing models for the first time. We show that the Bessel strategy is optimal for axisymmetric beams, while the Fourier approach can be applied to more general scenarios. Our work enables the efficient computation of a point-spread function on CPU or GPU, which can then be included in simulation and optimization pipelines.
Keywords point-spread function $\cdot$ localization microscopy $\cdot$ vectorial field propagation $\cdot$ open-source library
footnotetext: Abbreviations: PSF, point-spread function; NA, numerical aperture; FFT, fast Fourier transform; CPU, central processing unit; GPU, graphics processing unit; SMLM, single-molecule localization microscopy.
@@ 1 Introduction
The point-spread function (PSF), also referred to as the impulse response, is a fundamental concept that encapsulates key features of an optical microscope. Its monitoring and assessment have been longstanding routines in microscopy, supported by the development of specific tools [1, 2, 3] and the creation of a consortium to standardize best practices [4]. A detailed characterization of the PSF is essential to the design of computational imaging processes such as single-molecule localization microscopy [5] and super-resolution microscopy, including 3D deconvolution microscopy [6, 7], structured illumination microscopy (SR-SIM) [8, 9], fluctuation microscopy [10, 11], stimulated emission-depletion microscopy (STED) [12], and MINFLUX [13]. The PSF is at the heart of these localization techniques and helps one to achieve a spatial resolution beyond the optical diffraction limit.
@@
The PSF can be measured experimentally from a z-stack of tiny fluorescent beads [6, 14]. However, a theoretical PSF is often required. Here, the challenge lies in the accurate model of light propagation for imaging systems with high numerical aperture (NA). Detailed models have been proposed [15, 16, 17, 18] to (a) go beyond simple paraxial approximations, (b) take into account the vectorial nature of the electric field, and (c) include various aberration factors such as the Gibson-Lanni aberrations due to refraction at planar interfaces [19]. Several software packages have been developed to generate precise theoretical 3D PSFs that include various features and aberrations. Among these, the Huygens PSF softwarehttps://svi.nl/ is well-known in the microscopy community, and our Java-based ImageJ/FIJI plugin, PSF Generator [20], has also been widely used in the past. Other open-source alternatives include MATLAB-based solutions [21, 3, 22] as well as Python-based tools [23, 24].
@@
Computational PSF models have often been applied in fluorescence microscopy. For example, the fitting of a theoretical PSF to a measured one can improve accuracy [20, 25]. Detailed models enable the quantification of the 3D uncertainty [26], while the engineering of PSFs may take full advantage of parametric models [27, 28, 29]. They are a key component of high-quality single-molecule localization microscopy (SMLM) images [30, 31], including single-particle tracking [32, 33] and virtual SMLM microscopes [34, 35].
@@
Existing approaches to the computation of the PSF can be categorized into two classes of models: those based on 2D Fourier transforms [16, 3] or 1D integrals of Bessel functions [15, 36, 17, 18]. These two classes rely on different assumptions (with Bessel models being more restrictive) and bring different computational trade-offs. Most works focus on one of the two approaches, and the relationship between them remains unclear. For example, the Gibson-Lanni aberrations [19] or the apodization factor [15] are only applied in the Bessel case, although they could be generalized to both. To the best of our knowledge, no systematic benchmark of the two strategies has been published, either in terms of accuracy or computational speed.
@@
In this work, we propose a unifying framework for PSF models in which we show the equivalence between the Fourier and the Bessel approaches as being two separate parameterizations of the same propagation integral. This enables us to generalize diverse correction factors and apply them to both models. Eventually, the choice is simplified to scalar versus vectorial models, optionally with additional correction factors. We distribute an open-source PyTorch-based library called psf-generator. It inherits all the functionalities of PyTorch and allows for a seamless integration with modern learning-based algorithms. We benchmark the two classes of models on CPU and GPU and show that the Bessel model is advantageous for axisymmetric PSFs while the Fourier one is more general and thus more suitable for applications such as PSF engineering. To support the open-source community and facilitate the adoption of high-NA PSF models, we also provide a graphical interface in the napari ecosystem [37] and integrate our work in the chromatix optical simulation framework [38].
@@ 2 Background
Electromagnetic waves in optical systems are fundamentally described by the Helmholtz wave equations, yet solving these equations in full generality is computationally intractable. Various approximations and integral formulations have thus been developed. The cornerstone for precise PSF models in high-NA systems is the Richards-Wolf integral [15], which can be viewed as a vectorial extension of the Debye integral [39]. The conditions of validity for the Richards-Wolf integral have been thoroughly discussed by [40]. While alternative light propagation models exist, the Huygens-Fresnel approach has been shown to be equivalent, to some extent, for PSF calculations [41].
@@ 3 Theory
Figure 1: Geometry of the focusing of optical fields. An incident field $\mathbf{e}_{\mathrm{inc}}$ is transformed by a focusing element into a converging spherical wave $\mathbf{e}_{\infty}$. These fields are parameterized by a unit vector $\mathbf{s}$, either in Cartesian coordinates $(s_{x},s_{y},s_{z})$ or spherical coordinates $(\theta,\phi)$. The focus field $\mathbf{E}$ is parameterized by $\boldsymbol{\rho}=(x,y,z)$, rewritten here to introduce cylindrical coordinates $\rho$ and $\phi$.
@@ 3.1 The Richards-Wolf Model
As depicted in Figure 1, one obtains the PSF by computing the propagation of light after going through a focusing element, typically a microscope objective or lens. The incident field $\mathbf{e}_{\mathrm{inc}}(\mathbf{s})$, also called the pupil function, is represented by a disk with a maximal cutoff angle defined by the NA of the imaging system. The focusing element transforms the planar incident field into a spherical wave, $\mathbf{e}_{\infty}(\mathbf{s})$, evaluated on the Gaussian reference sphere. This corresponds to an ensemble of far fields that propagate with direction $\mathbf{s}$ and converge to the focal point $O$. Our goal is to compute the focused electric field $\mathbf{E}(\boldsymbol{\rho})$ near $O$. Due to the reciprocity of light propagation, this model can also be extended to the emission of a point source to the back focal plane of a microscope objective.
The model introduced by [15] is the starting point that allows us to derive all the precise PSF models described in previous works. The focal field is given by a sum of plane waves with direction $\mathbf{s}=(s_{x},s_{y},s_{z})$, as expressed by
$\mathbf{E}(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{\Omega}\mathbf{e}_{\infty}(\mathbf{s})\exp{\left(\mathrm{i}k\mathbf{s}\cdot\boldsymbol{\rho}\right)}\mathrm{d}\Omega,$ (1)
where $\boldsymbol{\rho}=(x,y,z)=(\rho\cos\varphi,\rho\sin\varphi,z)$ is the position vector in the focal region of the lens, $\mathbf{s}=(s_{x},s_{y},s_{z})=(\sin\theta\cos\phi,\sin\theta\sin\phi,\cos\theta)$ is a unit vector that describes the direction of an incoming ray, $f$ is the focal length of the lens, $k=\frac{2\pi n}{\lambda}$ is the wavenumber, $\lambda$ is the wavelength, $n$ is the refractive index of the propagation medium, and $\mathbf{e}_{\infty}(\mathbf{s})$ describes the field distribution on the Gaussian reference sphere. We integrate over the set $\Omega$ of solid angles defined on a region $s_{x}^{2}+s_{y}^{2}\leq s_{\mathrm{max}}^{2}$, where $s_{\mathrm{max}}=\frac{\mathrm{NA}}{n_{i}}$ is the cut-off determined by the NA and $n_{i}$ is the refractive index of the immersion medium. The angle $\theta$ is therefore defined within the immersion medium. We first describe the different classes of models based on this simplified concise equation and postpone to Section 3.4 the introduction of correction factors.
@@ 3.2 Scalar Models
As a first step, it is common to rely on a scalar approximation to simplify calculations, especially in low-NA scenarios. In this case, the far field is equal to the incident field, so that $e_{\infty}(\mathbf{s})=e_{\mathrm{inc}}(\mathbf{s})$. The focal field is then given by
$E(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{\Omega}e_{\infty}(s_{x},s_{y})\exp{\left(\mathrm{i}k\mathbf{s}\cdot\boldsymbol{\rho}\right)}\mathrm{d}\Omega.$ (2)
This expression involves a two-dimensional integral over the pupil disk. Two specific parameterizations yield the two classes of models described previously.
The Cartesian parameterization utilizes both $s_{x}$ and $s_{y}$ coordinates with $\mathrm{d}\Omega=\mathrm{d}s_{x}\mathrm{d}s_{y}/s_{z}$, which results in
$E(\boldsymbol{\rho})=-\frac{\mathrm{i}fk}{2\pi}\iint\limits_{s_{x}^{2}+s_{y}^{2}\leq s_{\max}^{2}}\frac{e_{\infty}(s_{x},s_{y})}{s_{z}}\exp{\left(\mathrm{i}ks_{z}z\right)}\exp{\left(\mathrm{i}k(s_{x}x+s_{y}y)\right)}\mathrm{d}s_{x}\mathrm{d}s_{y}.$ (3)
In this form, the focused field at a $z$ transverse plane is given by the 2D inverse Fourier transform of $(e_{\infty}(s_{x},s_{y})\exp{\left(\mathrm{i}ks_{z}z\right)}/s_{z})$, where $s_{z}=\sqrt{1-s_{x}^{2}-s_{y}^{2}}$. Thus, the Cartesian parameterization of the Richards-Wolf integral leverages the speed and efficiency of the fast Fourier transform (FFT) algorithm.
Alternatively, the spherical approach parameterizes the problem with two angles $\theta\in[0,\theta_{\max}]$ (the maximum angle $\theta_{\max}$ is determined by the NA) and $\phi\in[0,2\pi]$, as depicted in Figure 1. With $\boldsymbol{\rho}=(x,y,z)=(\rho\cos\varphi,\rho\sin\varphi,z)$ and $\mathrm{d}\Omega=\sin\theta\mathrm{d}\theta\mathrm{d}\phi$, the field in the focal region can be rewritten as
@@
$\exp{\left(\mathrm{i}k\rho\sin\theta\cos(\phi-\varphi)\right)}\exp{\left(\mathrm{i}kz\cos\theta\right)}\sin\theta.$ (4)
Equation (3.2) can be further simplified if one assumes that the pupil function is axisymmetric (rotational invariant), in the sense that $e_{\infty}(\theta,\phi)=e_{\infty}(\theta)$. In this case, the integral over $\phi$ in (3.2) can be computed explicitly using the Bessel function $J_{0}$The spherical parameterization often uses the following identities, where $J_{n}$ is the Bessel function of $n$th-order of the first kind: $\int_{0}^{2\pi}\cos(n\phi)\exp{\left(\mathrm{i}x\cos(\phi-\varphi)\right)}\mathrm{d}\phi=2\pi(\mathrm{i})^{n}J_{n}(x)\cos(n\varphi)$ $\int_{0}^{2\pi}\sin(n\phi)\exp{\left(\mathrm{i}x\cos(\phi-\varphi)\right)}\mathrm{d}\phi=2\pi(\mathrm{i})^{n}J_{n}(x)\sin(n\varphi)$ , which leads to:
$E(\boldsymbol{\rho})=-\mathrm{i}fk\int_{0}^{\theta_{\max}}\mathrm{d}\theta\,e_{\infty}(\theta)\,J_{0}(k\rho\sin\theta)\exp{\left(\mathrm{i}kz\cos\theta\right)}\sin\theta.$ (5)
Defocus is included in these models via the defocus phase factor $\exp{\left\{\mathrm{i}ks_{z}z\right\}}=\exp{\left(\mathrm{i}kz\cos\theta\right)}$ where $z$ is the defocus distance. This expression, also known as angular spectrum propagation [43], accurately models the propagation of an electric field in a homogeneous medium.